# SageAttention2 Technical Report: Accurate 4 Bit Attention for Plug-and-play Inference Acceleration

### v1

@@ Abstract
Although quantization for linear layers has been widely used, its application to accelerate the attention process remains limited. SageAttention utilizes 8-bit matrix multiplication, 16-bit matrix multiplication with 16-bit accumulator, and precision-enhancing methods, implementing an accurate and 2x speedup kernel compared to FlashAttention2. To further enhance the efficiency of attention computation while maintaining precision, we propose SageAttention2, which utilizes significantly faster 4-bit matrix multiplication (Matmul) alongside additional precision-enhancing techniques. First, we propose to quantize matrixes $(Q,K)$ to INT4 in a warp-level granularity and quantize matrixes $(\widetilde{P},V)$ to FP8. Second, we propose a method to smooth $Q$ and $V$, enhancing the accuracy of attention with INT4 $QK$ and FP8 $PV$. Third, we analyze the quantization accuracy across timesteps and layers, then propose an adaptive quantization method to ensure the end-to-end metrics over various models. The operations per second (OPS) of SageAttention2 surpass FlashAttention2 and xformers by about 3x and 5x on RTX4090, respectively. Comprehensive experiments confirm that our approach incurs negligible end-to-end metrics loss across diverse models—including those for large language processing, image generation, and video generation. The codes are available at https://github.com/thu-ml/SageAttention.
@@
Figure 1: The left figure shows the kernel speedup on RTX4090 GPU. The right figure shows the end-to-end inference speedup of generating the first token and performance metrics for the needle-in-a-haystack task (gkamradt, 2023) with a sequence length of 100K on Llama3.1 on L20 GPU.
@@ 1 Introduction
Due to the quadratic complexity of attention, efficient implementation becomes increasingly crucial as sequences lengthen in real-world applications (Jiang et al., 2024). Several strategies have been developed to mitigate computational demands of attention —such as (1) Linear attention methods (Wang et al., 2020; Choromanski et al., 2020; Yu et al., 2022; Katharopoulos et al., 2020) that reduce complexity to $O(N)$ and (2) Sparse attention methods (Liu et al., 2021; Chu et al., 2021; Li et al., ; Xiao et al., 2023b, 2024; Chen et al., 2023; Jiang et al., 2024; Venkataramanan et al., 2023; Gao et al., 2024; Fu et al., 2024) that selectively process parts of the context — these methods are only suitable for a limited range of models and tasks. Widely used attention methods optimize attention by exploiting hardware capacities to enhance the computation speed, such as FlashAttention (Dao et al., 2022), FlashAttention2 (Dao, 2023), FlashAttention3 (Shah et al., 2024), xformers (Lefaudeux et al., 2022), and SageAttention (Zhang et al., 2024). Those works do not omit the computation of attention over parts of the context and achieve impressive speed and precision performance across various models and tasks.
@@
For the two matrix multiplication (Matmul) operations in attention: $QK^{\top}$ and $PV$, SageAttention accelerates them by quantizing the $QK^{\top}$ to INT8 and uses FP16 Matmul with FP16 accumulators for $PV$. Moreover, to keep the accuracy of attention, SageAttention proposes smoothing $K$ by eliminating its channel-wise outliers. SageAttention achieves 2 $\times$ and 2.7 $\times$ speedup than FlashAttention2 and xformers and is the first quantized attention that incurs negligible end-to-end metrics loss across language, image, and video generation models. However, SageAttention has two weaknesses. (W1) INT8 Matmul achieves only half the speed of INT4. (W2) FP16 Matmul with FP16 accumulators is only compatible with RTX4090 and RTX3090 GPUs. To leverage the faster INT4 tensor cores for $QK^{\top}$ and using a method that can generally accelerate $PV$, we propose SageAttention2, which quantizes $Q,K$ to INT4 and $P,V$ to FP8.
@@
Challenges. Quantizing $Q,K$ to INT4 and $P,V$ to FP8 presents significant challenges. For example, when only per-tensor quantizing $Q,K$ to INT4, the text-to-video model CogvideoX will generate a completely blurry video (See Figure 2), and Llama2 only achieves a random-guessing-level accuracy of 25% on MMLU. After investigating deeply, we identified three primary challenges: (C1) The numerical range of INT4, which in quantization typically spans from $-7$ to $7$ and totals 15 numbers (Lin et al., 2024), leads to significant quantization errors when $Q$ and $K$ have some abnormal values. (C2) In specific layers and timesteps (for text-to-image/video) of some models, quantizing $Q$ and $K$ to INT4 and quantizing $P$ and $V$ to FP8 introduces noticeable errors in attention computation. These errors in worst-case layers/timesteps significantly impact the precision of the end-to-end output. (C3) We discover that the FP32 accumulator designed for FP8 matrix multiplication in the tensor core (mma.f32.f8.f8.f32) is actually FP22, specifically with 1 sign bit, 8 exponent bits, and 13 mantissa bits. This will lead to an accuracy loss of $PV$.
@@
Our approach. To address these challenges, we detailedly analyze the reasons and propose two methods. First, for the significant channel-wise outliers in matrices $Q$ and $K$, we adopt smoothing $K$ in SageAttention and propose an effective method to remove these outliers in $Q$. Specifically, we propose subtracting the average value of the channel dimension of $Q$, referred to as $\overrightarrow{Q}_{m}$. Then, we add $\overrightarrow{Q}_{m}K$ after the $QK$ Matmul to ensure the correctness of the attention computation. Second, we observe that specific layers and timesteps consistently present quantization challenges across different inputs. To maintain accuracy, we apply an adaptive mix precision method. Specifically, we employ 8-bit (INT8+FP8) attention for these problematic layers and timesteps and 4-bit (INT4+FP8) attention for the others. Third, to mitigate the accuracy loss associated with using the 22-bit accumulator for FP8 Matmul of $PV$, we propose a method to smooth $V$ to enhance the accuracy performance.
Performance. Importantly, we offer a high-performance implementation of SageAttention2 on RTX4090 and L20 GPUs. This implementation achieves a peak performance of 485 TOPS on the RTX4090, outperforming FlashAttention2 and xformers by approximately 3.1x and 5.4x, respectively. We extensively evaluate the end-to-end metrics of state-of-the-art text, image, and video generation models using SageAttention2. Across all models and tasks, SageAttention2 can be directly adopted in a plug-and-play manner with negligible loss in model performance.
@@ 2.1 FlashAttention
The computation of self-attention can be formulated as follows: $S=QK^{\top}/\sqrt{d},~{}P=\sigma(S),~{}O=PV$, where $\sigma(S)_{ij}=\exp(S_{ij})/\sum_{k}\exp(S_{ik})$ is the softmax operation. The matrices $Q$, $K$, and $V$ each have dimensions $N\times d$, while the matrix $S$, $P$ are $N\times N$. While $d$ is typically small, e.g., 64 or 128, $N$ can be thousands if not millions. Therefore, the $N\times N$ matrices $(S,P)$ are much larger than $(Q,K,V)$, and a naive implementation suffers from the huge amount of global memory IO for $(S,P)$ reads/writes. FlashAttention (Dao, 2023) proposes to tile $Q$, $K$, and $V$ from the token dimension into blocks $\{Q_{i}\},\{K_{i}\},\{V_{i}\}$ with block size of $b_{q}$, $b_{kv}$, $b_{kv}$, respectively. Then, to avoid the memory IO for $(S,P)$, it uses online softmax (Milakov & Gimelshein, 2018) to progressively compute each block of $O$, i.e., $O_{i}$:
@@
$S^{j}_{i}=Q_{i}K_{j}^{\top}/\sqrt{d},~{}~{}(m^{j}_{i},\widetilde{P}^{j}_{i})=\tilde{\sigma}(m^{j-1}_{i},S^{j}_{i}),~{}~{}l_{i}^{j}=$ $\exp(m^{j}_{i}-m^{j-1}_{i})l_{i}^{j-1}+\mathrm{rowsum}(\widetilde{P}^{j}_{i}),$
$O^{j}_{i}=\mathrm{diag}\left(\exp(m^{j}_{i}-m^{j-1}_{i})\right)O^{j-1}_{i}$ $+\widetilde{P}^{j}_{i}V_{j}$
@@ 2.2 Quantization
$\psi$ is a quantizer which converts a high-precision (e.g., FP32 or FP16) matrix $A$ to a low-precision format $\hat{A}$ (e.g., INT4 or FP8) with a scale $\delta_{A}$, and $\psi^{-1}$ is a dequantizer to convert back to high-precision. We should have $\psi^{-1}_{\delta_{A}}(\hat{A})\approx A$. The actual matrix multiplication $\hat{A}\hat{B}$ is carried in low-precision. In modern GPUs, low-precision matrix multiplication is usually multiple times faster than higher-precision ones. Many quantizers depend on the numerical format and granularity, e.g., how many elements share a common scale factor. For example, an INT4 per-tensor quantizer first computes the scale as the maximum absolute value of the entire tensor, scales the elements to the maximum representable range of INT4 [-7, +7], and then casts to INT4 with rounding: $\hat{A}=\lceil A/\delta_{A}\rfloor,\delta_{A}=\max(|{A}|)/7$. Likewise, per-token quantizer assigns a scale factor for each token of a tensor: $\hat{A}[i,:]=\lceil A[i,:]/\delta_{A}[i]\rfloor,\delta_{A}[i]=\max(|{A[i,:]}|)/7$. Also, per-channel quantizer assigns a scale factor for each channel of the tensor, i.e., along the channel dimension: $A[:,i]=\lceil A[:,i]/\delta_{A}[i]\rfloor,\delta_{A}[i]=\max(|A[:,i]|)/{7}$. Dequantization process is a element-wise scaling: $\psi_{\delta_{A}\delta_{B}}^{-1}(\hat{A}\hat{B})=\hat{A}\hat{B}*\delta_{A}*\delta_{B}$.
@@ 2.3 SageAttention
Where $Q_{i},K_{j}$ are the tilled block in FlashAttention, and $\text{mean}(K)$ is the average value of the channel dimension of $K$. SageAttention keeps $\widetilde{P}V$ as FP16 and uses FP16 Matmul with an FP16 accumulator for $\widetilde{P}V$. However, FP16 Matmul with an FP16 accumulator only has a speedup effect on RTX4090 and RTX3090 GPUs.
@@ 3 SageAttention-2
Figure 3: Workflow of SageAttention2. 1 Smooth Q,K,V. 2 A GEMV to obtain $\Delta S$. 3 Per-warp quantize Q,K and per-channel quantize V. 4 Perform the SageAttention2 kernel. 5 Correct the output.
@@ 3.1 Formulation
Table 1: End-to-end metrics comparison of different quantization methods, where Q,K are quantized into INT4, while P,V stay in full precision.
Q, K Smoothing (Q+K) Llama3.1 (Lambda) $\uparrow$ Llama3.1 (WikeText) $\downarrow$ CogVideo (vqa-a) $\uparrow$ CogVideo (vqa-t) $\uparrow$
@@
Table 2: Average accuracy across all layers using different quantization granularities.
Method Cos Sim $\uparrow$ Relative L1 $\downarrow$ RMSE $\downarrow$
Per-token 99.45% 0.0649 0.0335
Per-warp 99.45% 0.0648 0.0334
Per-block 98.03% 0.1492 0.0744
Per-tensor 97.15% 0.1800 0.0865
Table 3: Worst accuracy across all layers using different quantization granularities.
Method Cos Sim $\uparrow$ Relative L1 $\downarrow$ RMSE $\downarrow$
Per-token 96.76% 0.1916 0.0775
Per-warp 96.71% 0.1956 0.0779
Per-block 90.68% 0.3615 0.1490
Per-tensor 85.85% 0.4687 0.2261
3.2 Per-warp INT4 Quantization
SageAttention uses per-block quantization, which quantizes each block of $Q$ and $K$ for a GPU streaming processor. Such a quantization strategy could achieve an accuracy performance close to per-token quantization and avoid the overhead of the dot product of the quantization scale vectors $\delta_{Q}$ and $\delta_{K}$. However, quantizing $Q$ and $K$ to INT4 demands a more accurate quantization granularity. We propose per-warp quantization, a more precise and granular quantization approach than the per-block quantizer, also with no additional overhead of vectors dot product.
Specifically, each block of $Q$ in SageAttention will be split into $c_{w}$ segments, each processed by one of $c_{w}$ GPU warps in a GPU streaming processor (SM). Subsequently, each warp performs Matmul for its assigned segment. Per-warp INT4 quantization assigns a scale factor for every $b_{q}/{wc}$ tokens for $Q$:
$\hat{Q}[\frac{i*b_{q}}{c_{w}}$ $:\frac{b_{q}*(i+1)}{c_{w}},:]=\left\lceil\frac{Q[\frac{i*b_{q}}{c_{w}}:\frac{b_{q}*(i+1)}{c_{w}},:]}{\delta_{Q}[i]}\right\rfloor$ (4)
$\delta_{Q}[i]=\frac{\max(\left|\,Q[\frac{i*b_{q}}{c_{w}}:\frac{b_{q}*(i+1)}{c_{w}},:]\,\right|)}{7}$ (5)
Per-warp quantization will provide an accuracy performance superior to per-block quantization used in SageAttention, which we will discuss in Section 3.5.
@@ 3.3 Smooth Q
$S=QK=$ $\gamma(Q)\gamma(K)+\overrightarrow{Q_{m}}\gamma(K)$
@@
Where $\mathrm{mean}(K)$ and $\gamma(K)$ have been discussed in SageAttention (Zhang et al., 2024), and the transformation for $K$ does not change the attention score $\widetilde{P}$. $\mathrm{mean}(Q)=\frac{1}{N}\sum_{t=1}^{N}Q[t,:]$ is a vector with a shape $1\times d$. For the transformation of $Q$, we will perform a GEMV (general matrix-vector multiplication) between $\overrightarrow{Q_{m}}$ and $K^{\top}$, i.e., $\overrightarrow{Q_{m}}K^{\top}$ as $\Delta S$. This $\Delta S$ will be added to $S$ during the attention computation. Finally, the transformation from full-precision $Q,K$ to quantized $\hat{Q},\hat{K}$ can be expressed as shown in Eqution 6, where $\psi_{Q}$, $\psi_{K}$ is two quantizers for $Q$ and $K$. In other words, full-precision $Q,K$ are substracted with the means of the channel dimension before being quantized.

### v3

# SageAttention2: Efficient Attention with Thorough Outlier Smoothing and Per-thread INT4 Quantization

@@ Abstract
Although quantization for linear layers has been widely used, its application to accelerate the attention process remains limited. To further enhance the efficiency of attention computation compared to SageAttention while maintaining precision, we propose SageAttention2, which utilizes significantly faster 4-bit matrix multiplication (Matmul) alongside additional precision-enhancing techniques. First, we propose to quantize matrixes $(Q,K)$ to INT4 in a hardware-friendly thread-level granularity and quantize matrixes $(\widetilde{P},V)$ to FP8. Second, we propose a method to smooth $Q$, enhancing the accuracy of INT4 $QK^{\top}$. Third, we propose to use an FP32 Matmul buffer for $PV$ to enhance the accuracy of FP8 $\widetilde{P}V$. The operations per second (OPS) of SageAttention2 surpass FlashAttention2 and xformers by about 3x and 5x on RTX4090, respectively. Comprehensive experiments confirm that our approach incurs negligible end-to-end metrics loss across diverse models—including those for large language processing, image generation, and video generation. The codes are available at https://github.com/thu-ml/SageAttention.
@@
Figure 1: The upper left figure shows the kernel speedup on RTX4090 GPU. The upper right figure shows the end-to-end inference speedup of generating the first token and performance metrics for the needle-in-a-haystack task (gkamradt, 2023) with a sequence length of 100K on Llama3.1 on L20 GPU. The figure below shows two videos generated by CogvideoX (1.5-5B) using FlashAttention2 and SageAttention2 on RTX4090, showing that SageAttention2 accelerates generation 1.8x with no video quality loss.
@@ 1 Introduction
Due to the quadratic time complexity of attention, efficiently implementing it becomes crucial as sequence lengths increase in real-world applications (Jiang et al., 2024). Several strategies have been developed to mitigate computational demands of attention —such as (1) Linear attention methods (Wang et al., 2020; Choromanski et al., 2020; Yu et al., 2022; Katharopoulos et al., 2020) that reduce complexity to $O(N)$ and (2) Sparse attention methods (Liu et al., 2021; Chu et al., 2021; Li et al., ; Xiao et al., 2023b, 2024; Chen et al., 2023; Jiang et al., 2024; Venkataramanan et al., 2023; Gao et al., 2024; Fu et al., 2024) that selectively process parts of the context — these methods are only suitable for a limited range of models and tasks. Widely used attention methods optimize attention by exploiting hardware capacities to enhance the computation speed, such as FlashAttention (Dao et al., 2022), FlashAttention2 (Dao, 2023), FlashAttention3 (Shah et al., 2024), xformers (Lefaudeux et al., 2022), and SageAttention (Zhang et al., 2024). Those works do not omit the computation of attention over parts of the context and achieve impressive speed and precision performance across various models and tasks.
@@
Motivation. For the two matrix multiplication (Matmul) operations in attention: $QK^{\top}$ and $\widetilde{P}V$, SageAttention accelerates them by quantizing the $QK^{\top}$ to INT8 and uses FP16 Matmul with FP16 accumulators for $\widetilde{P}V$. Moreover, to keep the accuracy of attention, SageAttention proposes smoothing $K$ by eliminating its channel-wise outliers. SageAttention achieves 2 $\times$ and 2.7 $\times$ speedup than FlashAttention2 and xformers and is the first quantized attention that incurs negligible end-to-end metrics loss across language, image, and video generation models. However, SageAttention has two weaknesses. (W1) INT8 Matmul achieves only half the speed of INT4. (W2) FP16 Matmul with FP16 accumulators is only compatible with RTX4090 and RTX3090 GPUs. To leverage the faster INT4 tensor cores for $QK^{\top}$ and using a method that can accelerate $\widetilde{P}V$ on a broader range of GPUs, we propose SageAttention2, which quantizes $Q,K$ to INT4 and $\widetilde{P},V$ to FP8.
@@
Challenges. Quantizing $Q,K$ to INT4 and $\widetilde{P},V$ to FP8 presents significant challenges. For example, when only per-tensor quantizing $Q,K$ to INT4, the text-to-video model CogvideoX will generate a completely blurry video (See Figure 2), and Llama2 only achieves a random-guessing-level accuracy of 25% on MMLU. After investigating deeply, we identified three primary challenges: (C1) The numerical range of INT4 is quite limited compared to FP16 or INT8, leading to significant quantization errors when $Q$ and $K$ have some abnormal values. (C2) We discover that the FP32 accumulator designed for FP8 matrix multiplication in the tensor core (mma.f32.f8.f8.f32) is actually FP22, specifically with 1 sign bit, 8 exponent bits, and 13 mantissa bits. This will lead to an accuracy loss of $\widetilde{P}V$.
@@
Our approach. To address (C1), we first discovered that the per-block quantization of $Q,K$ in SageAttention does not achieve sufficient accuracy for INT4 quantization. Simultaneously, to avoid the extra latency caused by per-token dequantization, where each GPU thread corresponds to multiple quantization scales, we propose an adequately precise quantization method that incurs no additional dequantization overhead. Specifically, we introduce a per-thread quantization approach based on the mapping between the GPU thread and memory layout of matrices as dictated by the PTX mma instructions. This method groups tokens corresponding to the same thread for quantization and dequantization, ensuring that each thread is associated with a single quantization scale. This approach achieves much better accuracy performance than per-block quantization with no additional dequantization overhead. Second, for the significant channel-wise outliers in matrices $Q$ and $K$, we adopt smoothing $K$ in SageAttention and propose an effective method to remove these outliers in $Q$. Specifically, we propose subtracting the average value of the channel dimension of $Q$, referred to as $\overrightarrow{Q}_{m}$. Then, we add $\overrightarrow{Q}_{m}K$ after the $QK^{\top}$ Matmul to ensure the correctness of the attention computation. To address (C2)—the accuracy loss from using a 22-bit accumulator for FP8 Matmul of $\widetilde{P}V$—, we propose to use an FP32 buffer to accumulate the values from the 22-bit accumulator after each block Matmul of $\widetilde{P}V$. This method confines the errors to the block range. Additionally, we design an optional technique to enhance the accuracy of the 22-bit accumulator. Specifically, we could smooth $V$ by subtracting the average value of its channel dimension and adding the subtracted item to the attention output to maintain the correctness.
Performance. Importantly, we offer a high-performance implementation of SageAttention2 on RTX4090 and L20 GPUs. This implementation achieves a peak performance of 481 TOPS on the RTX4090, outperforming FlashAttention2 and xformers by approximately 3x and 5x, respectively. Note that FlashAttention3 is tailored to and can only be used with the Nvidia Hopper architecture, i.e., H100 and H800. We extensively evaluate the end-to-end metrics of state-of-the-art text, image, and video generation models using SageAttention2. Across all models and tasks, SageAttention2 can be directly adopted in a plug-and-play manner with negligible loss in model performance.
@@ 2.1 FlashAttention
The computation of self-attention can be formulated as follows: $S=QK^{\top}/\sqrt{d},~{}P=\sigma(S),~{}O=PV$, where $\sigma(S)_{ij}=\exp(S_{ij})/\sum_{k}\exp(S_{ik})$ is the softmax operation. The matrices $Q$, $K$, and $V$ each have dimensions $N\times d$, while the matrix $S$, $P$ are $N\times N$. While $d$ is typically small, e.g., 64 or 128, $N$ can be millions. Therefore, the $N\times N$ matrices $(S,P)$ are much larger than $(Q,K,V)$, and a naive implementation suffers from the huge amount of global memory IO for $(S,P)$ reads/writes. FlashAttention (Dao, 2023) proposes to tile $Q$, $K$, and $V$ from the token dimension into blocks $\{Q_{i}\},\{K_{i}\},\{V_{i}\}$ with block size of $b_{q}$, $b_{k}$, $b_{v}$, respectively, where $b_{k}$ = $b_{v}$. Then, to avoid the memory IO for $(S,P)$, it uses online softmax (Milakov & Gimelshein, 2018) to progressively compute each block of $O$, i.e., $O_{i}$:
@@
$S^{j}_{i}=Q_{i}K_{j}^{\top}/\sqrt{d},~{}~{}(m^{j}_{i},\widetilde{P}^{j}_{i})=\tilde{\sigma}(m^{j-1}_{i},S^{j}_{i}),~{}~{}l_{i}^{j}=$ $\exp(m^{j-1}_{i}-m^{j}_{i})l_{i}^{j-1}+\mathrm{rowsum}(\widetilde{P}^{j}_{i}),$
$O^{j}_{i}=\mathrm{diag}\left(\exp(m^{j-1}_{i}-m^{j}_{i})\right)O^{j-1}_{i}$ $+\widetilde{P}^{j}_{i}V_{j}$
@@ 2.2 Quantization
$\psi$ is a quantizer which converts a high-precision (e.g., FP32 or FP16) matrix $A$ to a low-precision format $\hat{A}$ (e.g., INT4 or FP8) with a scale $\delta_{A}$, and $\psi^{-1}$ is a dequantizer to convert back to high-precision. We should have $\psi^{-1}_{\delta_{A}}(\hat{A})\approx A$. The actual matrix multiplication $\hat{A}\hat{B}$ is carried in low precision. In modern GPUs, low-precision matrix multiplication is usually multiple times faster than higher-precision ones. Many quantizers depend on the numerical format and granularity, e.g., how many elements share a common scale factor. For example, an INT4 per-tensor quantizer first computes the scale as the maximum absolute value of the entire tensor, scales the elements to the maximum representable range of INT4 [-7, +7], and then casts to INT4 with rounding: $\hat{A}=\lceil A/\delta_{A}\rfloor,\delta_{A}=\max(|{A}|)/7$. Likewise, per-token quantizer assigns a scale factor for each token of a tensor: $\hat{A}[i,:]=\lceil A[i,:]/\delta_{A}[i]\rfloor,\delta_{A}[i]=\max(|{A[i,:]}|)/7$. Also, per-channel quantizer assigns a scale factor for each channel of the tensor, i.e., along the channel dimension: $A[:,i]=\lceil A[:,i]/\delta_{A}[i]\rfloor,\delta_{A}[i]=\max(|A[:,i]|)/{7}$. The dequantization process is an element-wise scaling. For example, for per-tensor dequantization, $\psi_{\delta_{A}\delta_{B}}^{-1}(\hat{A}\hat{B})=\hat{A}\hat{B}*\delta_{A}*\delta_{B}$.
@@ 2.3 SageAttention
Where $Q_{i},K_{j}$ are the tilled block in FlashAttention, and $\text{mean}(K)$ is the average value of the channel dimension of $K$. Additionally, SageAttention keeps $\widetilde{P}V$ as FP16 and uses FP16 Matmul with an FP16 accumulator for $\widetilde{P}V$. However, FP16 Matmul with an FP16 accumulator only has a speedup effect on RTX4090 and RTX3090 GPUs.
@@ 3 SageAttention-2
Figure 3: Workflow of SageAttention2. 1 Smooth Q,K,V. 2 A GEMV to obtain $\Delta S$. 3 Per-thread quantize Q,K and per-channel quantize V. 4 Perform the SageAttention2 kernel. 5 Correct the output.
@@ 3.1 Formulation
Table 1: End-to-end metrics comparison of different quantization methods, where $Q,K$ are quantized into INT4, while $\widetilde{P},V$ stay in full precision.
Q, K Smoothing (Q+K) Llama3.1 (Lambda) $\uparrow$ Llama3.1 (WikeText) $\downarrow$ CogVideoX (vqa-a) $\uparrow$ CogVideoX (vqa-t) $\uparrow$
@@ 3.2 Smooth Q
$S=QK^{\top}=$ $\gamma(Q)\gamma(K)^{\top}+\overrightarrow{Q_{m}}\gamma(K)^{\top}$
@@
Where $\mathrm{mean}(K)$ and $\gamma(K)$ have been discussed in SageAttention (Zhang et al., 2024), and the transformation for $K$ does not change the attention score $\widetilde{P}$. $\mathrm{mean}(Q)=\frac{1}{N}\sum_{t=1}^{N}Q[t,:]$ is a vector with a shape $1\times d$. For the transformation of $Q$, we will perform a GEMV (general matrix-vector multiplication) between $\overrightarrow{Q_{m}}$ and $K^{\top}$, i.e., $\overrightarrow{Q_{m}}K^{\top}$ as $\Delta S$. This $\Delta S$ will be added to $S$ during the attention computation. Finally, the transformation from full-precision $Q,K$ to quantized $\hat{Q},\hat{K}$ can be expressed as shown in Eqution 4, where $\psi_{Q}$, $\psi_{K}$ is two quantizers for $Q$ and $K$. In other words, full-precision $Q,K$ are substracted with the means of the channel dimension before being quantized.
@@
Figure 6: An example of per-thread quantization. The left figure shows the correspondence between the quantization scales and the tokens in each GPU warp. The right figure shows the correspondence between quantization tokens and GPU threads in a MMA.m16n8k64 instruction, showing that each GPU thread only corresponds to one quantization scale in $\delta_{Q}$ and $\delta_{K}$ in dequantization.
Table 2: Average accuracy across all layers of CogvideoX using different quantization granularities.
Method Cos Sim $\uparrow$ Relative L1 $\downarrow$ RMSE $\downarrow$
Per-token 99.45% 0.0649 0.0335
Per-thread 99.45% 0.0622 0.0313
Per-block 98.03% 0.1492 0.0744
Per-tensor 97.15% 0.1800 0.0865
Table 3: Worst accuracy across all layers of CogvideoX using different quantization granularities.
Method Cos Sim $\uparrow$ Relative L1 $\downarrow$ RMSE $\downarrow$
Per-token 96.76% 0.1916 0.0775
Per-thread 96.72% 0.1932 0.0776
Per-block 90.68% 0.3615 0.1490
Per-tensor 85.85% 0.4687 0.2261
3.3 INT4 Per-thread Quantization
SageAttention uses per-block quantization, which quantizes each block of $Q$ and $K$ for a GPU streaming processor. Such a quantization strategy could achieve an accuracy performance close to per-token quantization and avoid the overhead of the dot product of the quantization scale vectors $\delta_{Q}$ and $\delta_{K}$. However, quantizing $Q$ and $K$ to INT4 demands a more accurate quantization granularity. We propose per-thread quantization, a more precise and granular quantization approach than the per-block quantizer, also without the additional overhead of vectors dot product required for per-token quantization.
Specifically, each block of $Q$, i.e., $Q_{i}$, in SageAttention will be split into $c_{w}$ segments and processed by $c_{w}$ GPU warps in a GPU streaming processor (SM). We call each segment of $Q_{i}$ as $Q_{w}$, and $k_{w}=K_{j}$ because a $K_{j}$ only corresponds to one warp. Then, each warp containing 32 threads uses the mma.m16n8k64 PTX instruction (NVIDIA, ) for the $Q_{w}K_{j}^{\top}$. According to the layout requirement of this instruction, we find that the $Q_{w}[8\times(n\%8)]$ could share one quantization scale, and $K_{j}[8\times(n\%8)]$ and $K_{j}[8\times(n\%8+1)]$ could share one quantization scale, where $n$ is the token index. Such a quantization method is more fine-grained and with no additional overhead. This is because it assigns different GPU threads to distinct quantization groups based on the MMA instruction layout, with each thread performing dequantization only using a single quantization scale value. The per-thread quantization process can be formulated as follows.
$i_{\delta q}=$ $\lfloor(n*8*c_{w}/b_{q})\rfloor$
$q_{i}[i_{\delta q}]=$$\{8\times(n\%8)+\lfloor(n*c_{w}/b_{q})\rfloor*b_{q}/c_{w}\},~{}~{}n\in[0,N]$
$\delta_{Q}[i_{\delta q}]=$ $\frac{\max(\left|\,Q[q_{i}[i_{\delta q}]]\,\right|)}{7}$
$\hat{Q}[q_{i}[i_{\delta q}]]=$ $\left\lceil\frac{Q[q_{i}[i_{\delta q}]]}{\delta_{Q}[i_{\delta q}]}\right\rfloor$ (5)
$i_{\delta k}=$ $\lfloor(n*4/b_{k})\rfloor$
$k_{n}[i_{\delta k}]=$$\{8\times(n\%8)+\lfloor n/b_{k}\rfloor*b_{k}\}\cup\{8\times(n\%8)+1+\lfloor n/b_{k}\rfloor*b_{k}\},~{}~{}n\in[0,N]$
$\delta_{K}[i_{\delta k}]=$ $\frac{\max(\left|\,K[k_{n}[i_{\delta k}]]\,\right|)}{7}$
$\hat{K}[k_{n}[i_{\delta k}]]=$ $\left\lceil\frac{K[k_{n}[i_{\delta k}]]}{\delta_{K}[i_{\delta k}]}\right\rfloor$ (6)
We show an example of per-thread quantization in Figure 6.